# EasyRL4Rec: An Easy-to-use Library for Reinforcement Learning Based Recommender Systems

### v2

@@ 1. Introduction
  • Distinct evaluation strategies in different work. The absence of standardized evaluation metrics complicates the comparison of model performances across different research teams. While some studies follow traditional RS metrics such as Normalized Discounted Cumulative Gain (NDCG) and Hit Rate (HR), others rely on RL-specific measures like cumulative reward and interaction length, leading to inconsistencies in performance assessment.
@@
To tackle the above issues, we implement a comprehensive code library for RL-based RSs, named EasyRL4Rec. It provides an easy-to-use framework with core modules with rich choices and a unified training and evaluating process, aiming to simplify the model development and experimental process in the domain of RL-based RS. The library is composed of four core modules: Environment, Policy, StateTracker, and Collector, which cater to different stages of the RL interaction process. Environment, built from lightweight static datasets, provides feedback on upcoming actions. The Policy module selects the optimal action based on the current state, which is encoded by StateTracker. Collector bridges the interactions between Environment and Policy. To address the difficulty of obtaining user states from environments, we implement multiple StateTrackers for state modeling, encompassing today’s popular methods in sequential modeling (Hidasi et al., 2015; Tang and Wang, 2018; Kang and McAuley, 2018; Yuan et al., 2019). Since items in recommendation systems are discrete, EasyRL4Rec includes a mechanism to convert continuous actions to discrete items, allowing for continuous action-based policies.
Moreover, EasyRL4Rec provides a unified training and evaluation procedure by Trainer and Evaluator executors. Evaluator supports three modes, allowing for the removal of recommended items and a quit mechanism. The Trainer module offers two training paradigms: learning directly from offline logs or training with a pre-trained user model. With this unified framework, we conduct comprehensive experiments on classic RL models and several recent work (Xin et al., 2020; Gao et al., 2023) and present insightful results. The main contributions of this work can be summarized as follows:
@@
  • Tailored Designs for Recommendation Scenarios. In response to challenges when applying RL algorithms in practical recommender systems, we have developed customizable modules for state modeling and action representation, with a conversion mechanism to support continuous action-based policies.
  • Insightful Experiments for RL-based RSs. With EasyRL4Rec, we conduct comprehensive experiments to compare the performance of classic RL models and some recent work. We present insightful experimental results from various perspectives.
@@ 2.1. Development of RL-Based RSs
Reinforcement Learning (RL), a branch of machine learning, focuses on agents learning the ability of decision-making through environmental feedback [^2] RL in this work specifically refers to deep reinforcement learning.. Recently, RL-Based Recommender Systems (RSs) have gained considerable attention due to their ability to model recommendation as a multi-step decision-making process and enhance the long-term benefits.
@@ 2.2. Resources for RL-Based RSs
Existing resources for RL-based RSs can be categorized into two groups: Simulators & Datasets, and Frameworks & Libraries. In Table  1, we summarize the characteristics of existing resources and our EasyRL4Rec regarding modules and experimental procedures.
@@ 4.1. Library Overview
The overall architecture of EasyRL4Rec is illustrated in Figure 1. The library is composed of four core modules: Environment, Policy, StateTracker, and Collector, each addressing distinct stages of the reinforcement learning interaction process. These modules build upon the foundations provided by the Gymnasium and Tianshou (Weng et al., 2022) libraries. The Buffer serves as a fundamental data structure for organizing raw data trajectories, while the Trainer and Evaluator act as executors, managing the entire process. In the following, we will delve into the designs of these core modules. The details of the training and evaluation pipeline can be found in Section 5.
@@ 4.3. Policy
The Policy module applies a reinforcement learning algorithm to select the optimal action based on the current state. We implement this module by extending RL policies in the Tianshou (Weng et al., 2022), incorporating the following tailored designs for recommendation scenarios.
Firstly, we support both discrete and continuous action-based policies. Given that items in RSs are discrete and more suitable for discrete action-based policies, EasyRL4Rec includes a mechanism to convert continuous actions to discrete items, supporting continuous action-based policies. Secondly, we customize policies by encoding the state via StateTracker since states cannot be directly obtained from environments. These encoded state embeddings are optimized simultaneously with policies. Thirdly, we introduce a Remove Recommended Items option when interacting with environments, addressing the common need for multi-round recommendation. This feature is implemented by adding a mask for logits.
@@ 4.5. Collector
Collector plays a pivotal role in both the Training and Evaluation stages. To be specific, at time $t$, the collector would call the Policy module to execute an action $a_{t}$ according to observed information $o_{t}$. Subsequently, it conveys this action to Environment and gets corresponding rewards $r_{t}$. Considering a complete interaction from time $1$ to time $T$, the observations, actions, and rewards at each timestamp, denoted as $\{(o_{1},a_{1},r_{1}),...,(o_{T},a_{T},r_{T})\}$, would be considered as one single trajectory and stored in Buffer. These trajectories would be sampled for policy learning and subsequent updates.
@@ 7. Experiments
Table 4. Performance comparison between model-free RL methods trained with a user model on three datasets. Policies using continuous action are indicated by the symbol (C). The meaning of underlining and bold should be included in caption. The best results are highlighted in bold, and the second-best results are underlined.
@@ 7.2. Performance of Model-free RL
Firstly, discrete action-based methods perform much better than continuous action-based methods (marked by (C)). Discrete methods such as DQN and C51 in the off-policy category, and PG and PPO in the on-policy category, achieve higher $\text{R}_{\text{cumu}}$ and longer $Length$ across all three datasets. Yet continuous methods like DDPG(C), TD3(C), PG(C), and A2C(C) show fewer Lengths, which might be due to less efficient exploration strategies or poor fit between the continuous action representation and the discrete item space in RSs. This trend suggests that discrete action spaces, which are more aligned with the discrete nature of items in recommender systems, are more effective for these model-free RL algorithms.

### v3

@@ 1. Introduction
  • Distinct evaluation strategies in different work. The absence of standardized evaluation metrics complicates model comparisons across different research teams. While some studies follow traditional RS metrics such as Normalized Discounted Cumulative Gain (NDCG) and Hit Rate (HR), others use RL-specific measures like cumulative reward and interaction length, leading to inconsistencies in performance assessment.
@@
To tackle the above issues, we implement a comprehensive code library for RL-based RSs, named EasyRL4Rec. It provides an easy-to-use framework with core modules with rich choices and a unified training and evaluating process, aiming to simplify the model development and experimental process in the domain of RL-based RS. The library is composed of four core modules: Environment, Policy, StateTracker, and Collector, which cater to different stages of the RL interaction process. Environment, built from lightweight static datasets, provides feedback on upcoming actions. The Policy module selects the optimal action based on the current state, which is encoded by StateTracker. Collector bridges the interactions between Environment and Policy. To address the difficulty of obtaining user states from environments, we implement multiple StateTrackers for state modeling, encompassing today’s popular methods in sequential modeling (Hidasi et al., 2015; Tang and Wang, 2018; Kang and McAuley, 2018; Yuan et al., 2019). Since items in RSs are discrete, EasyRL4Rec includes a mechanism to convert continuous actions to discrete items, allowing for continuous action-based policies.
Moreover, EasyRL4Rec provides a unified training and evaluation procedure by Trainer and Evaluator executors. Evaluator supports three modes, allowing for the removal of recommended items and a quit mechanism. Trainer offers two training paradigms: learning directly from offline logs or training with a pre-trained user model. With this unified framework, we conduct comprehensive experiments on classic RL models and several recent work (Xin et al., 2020; Gao et al., 2023) and present insightful results. The main contributions of this work can be summarized as follows:
@@
  • Tailored Designs for Recommendation Scenarios. In response to challenges when applying RL algorithms in practical RSs, we have developed customizable modules for state modeling and action representation, with a conversion mechanism to support continuous action-based policies.
  • Insightful Experiments for RL-based RSs. With EasyRL4Rec, we conduct experiments to compare the performance of classic RL models and some recent work, presenting insightful results from multiple perspectives.
@@ 2.1. Development of RL-Based RSs
Reinforcement Learning (RL), a branch of machine learning, focuses on agents learning the ability of decision-making through environmental feedback [^2] RL in this work specifically refers to deep reinforcement learning.. Recently, RL-Based Recommender Systems have gained considerable attention due to their ability to model recommendation as a multi-step decision-making process and enhance the long-term benefits.
@@ 2.2. Resources for RL-Based RSs
Existing resources for RL-based RSs can be categorized into two groups: Simulators & Datasets, and Frameworks & Libraries. In Table  1, we summarize the characteristics of existing resources regarding modules and experimental procedures.
@@ 4.1. Library Overview
The overall architecture of EasyRL4Rec is illustrated in Figure 1. The library is composed of four core modules: Environment, Policy, StateTracker, and Collector, each addressing distinct stages of the reinforcement learning interaction process. The Buffer serves as a fundamental data structure for organizing raw data trajectories, while the Trainer and Evaluator act as executors, managing the entire process. In the following, we will delve into the designs of these core modules. The details of the training and evaluation pipeline can be found in Section 5.
@@ 4.3. Policy
The Policy module applies a RL algorithm to select the optimal action based on the current state. We implement this module by extending RL policies in the Tianshou (Weng et al., 2022), incorporating the following tailored designs for recommendation scenarios.
Firstly, we support both discrete and continuous action-based policies. Given that items in RSs are discrete and better suited for discrete actions, we include a mechanism to convert continuous actions to discrete items, supporting continuous action-based policies. Secondly, we customize policies by encoding the state via StateTracker since states cannot be directly obtained from environments. These encoded state embeddings are optimized simultaneously with policies. Thirdly, we introduce a Remove Recommended Items option when interacting with environments, addressing the common need for multi-round recommendation. This feature is implemented by adding a mask for logits.
@@ 4.5. Collector
Collector plays a pivotal role in both the Training and Evaluation stages. To be specific, at time $t$, the Collector would call Policy to execute an action $a_{t}$ according to observed information $o_{t}$. Subsequently, it conveys this action to Environment and gets corresponding rewards $r_{t}$. Considering a complete interaction from time $1$ to time $T$, the observations, actions, and rewards at each timestamp, denoted as $\{(o_{1},a_{1},r_{1}),...,(o_{T},a_{T},r_{T})\}$, would be considered as one single trajectory and stored in Buffer. These trajectories would be sampled for policy learning and subsequent updates.
@@ 7. Experiments
Table 4. Performance comparison between model-free RL methods trained with a user model on three datasets. Policies using continuous action are indicated by the symbol (C). The meaning of underlining and bold should be included in the caption. The best results are highlighted in bold, and the second-best results are underlined.
@@ 7.2. Performance of Model-free RL
Firstly, discrete action-based methods perform much better than continuous action-based methods (marked by (C)). Discrete methods such as DQN and C51 in the off-policy category, and PG and PPO in the on-policy category, achieve higher $\text{R}_{\text{cumu}}$ and longer $Length$ across all three datasets. Yet continuous methods like DDPG(C), TD3(C), PG(C), and A2C(C) show fewer Lengths, which might be due to less efficient exploration strategies or poor fit between the continuous action representation and the discrete item space in RSs. This trend suggests that discrete action spaces, which are more aligned with the discrete items in recommender systems, are more effective for model-free RL algorithms.